OpenStudy (anonymous):

Apparently cos((9 * pi) / 4) = cos((pi / 4) + 2 * pi) Can someone explain why this is?

OpenStudy (agent0smith):

This is because 9pi/4 is equal to pi/4 + 2pi... try it by putting the 2pi over a common denominator: \[\large \frac{ \pi }{ 4} + 2 \pi = \frac{ \pi }{ 4} + \frac{ 8 \pi }{ 4} = ?\]

OpenStudy (anonymous):

yes...^^^ and also, 9pi/4 and pi/4 are coterminal angles.... that is, the terminal side is the same position for both angles.